Beginning three months from now, you want to be able to withdraw $4,900 each quarter from your bank account to cover college expenses. The account pays 1.87 percent interest per quarter. How much do you need to have in your account today to meet your expense needs over the next four years? A. $78,400.00 B. $71,086.02 C. $67,221.04 D. $48,785.91
Answer #1

Rate Per quarter =1.87%
PMT =4900
Number of Periods =4*4 =16
PV using annuity due formula =PMT*((1-(1+r)^-n)/r) =4900*((1-(1+1.87%)^-16)/1.87%) =67221.04

Option c is correct option